In a certain region of space, the electric potential is V (x, y, z) = Axy − Bx2 + Cy, where A, B, and
C are positive constants, (a) Calculate the x-, y- and z-components of the electric field.(b) At what
point is the electric field equal to zero?
1
Expert's answer
2021-07-04T17:45:22-0400

Gives

V(x,y,z)=Axy-Bx2-Cy

"E=-\nabla V"

"E_x=-\frac{dV}{dx}=-(Ay-2Bx)"

"E_y=-\frac{dV}{dy}=-(Ax+c)"

"E_z=-\frac{dV}{dz}=0"

"E=E_x+E_y+E_z"

Find out point where electric field is zero


"E=-\hat{i}(Ay-2Bx)-\hat{j}(Ax+c)+\hat{k}"

E=0

"\hat{i} components zero"

Ax+c=0

"x=-\frac{c}{A}"

Ay-2Bx=0

"y=\frac{2Bx}{A}"

Put x value

"y=-\frac{2Bc}{A^2}"

Point "(x,y,z)=(-\frac{c}{A},\frac{2Bc}{A^2},0)"